# Inexact Restoration via random models for unconstrained noisy optimization

### v2

@@ 4.3. Iteration complexity
We conclude our analysis by deriving the expected number $\mathbb{E}[{\mathcal{K}}_{\epsilon}]$ for the hitting time given in Definition 4.1. We rely on results for random processes introduced in [9, §2] and, to this end, we need some technical results that are proved in this section. In particular, we exploit the Lyapunov function:
@@
Here $v\in(0,1)$ is a prefixed constant to be later specified and $\Sigma$ is a constant such that
(4.14) ${f(x_{k})}-h(y_{k})+\Sigma\geq 0,\quad\forall\ k\geq 0.$
@@
(4.15) $\mathcal{G}_{k,1}$ =$=\left\{|f(X_{k})-\bar{F}_{k}^{\dagger}|\leq\kappa\min\{\Delta_{k}^{2},h(Y_{k})\}\right\},$
(4.16) $\mathcal{G}_{k,2}$ =$=\left\{|f(X_{k})-\bar{F}_{k}^{*}|\leq\kappa\min\{\Delta_{k}^{2},h(Y_{k})\}\right\},$
(4.17) $\mathcal{G}_{k,3}$ =$=\left\{\|\nabla f(X_{k})-G_{k}\|\leq\kappa\Delta_{k}\right\},$
(4.18) $\mathcal{G}_{k,4}$ =$=\left\{|f(X_{k}+P_{k})-\bar{F}_{k}^{p}|\leq\kappa\Delta_{k}^{2}\right\},$
@@
We note that iteration $k$ is true by Definition 4.3 when $\mathcal{I}_{k}\mathcal{J}_{k}=1$.
@@
Note that Assumption 4.9 is satisfied whenever the random variables $\omega^{\dagger}_{y_{k+1}^{t}}$, $\omega^{*}_{y_{k+1}^{t}}$ and $\omega_{y_{k+1}^{t}}^{g}$ that model the noise are independent.
@@
Suppose that Assumptions 2.1 and 4.9 hold. Given a probability $\pi\in(0,1]$, let $\bar{\mu}(\pi^{\frac{1}{4}})>0$ be defined as in Assumption 2.1(ii) and assume $\mu\leq\kappa\bar{\mu}(\pi^{\frac{1}{4}})$. Then we have
@@
Let $\kappa$, $\zeta$ and $v\in(0,1)$ be some scalars satisfying
@@
The following theorem shows a bound on the expected value $\mathbb{E}_{k-1}[\Psi_{k+1}-\Psi_{k}]$ from which one easily derives that $\Delta_{k}$ tends to zero almost surely.
@@
Let $\kappa,\,\zeta,\,C_{1},\,C_{2},\,C_{3},\,C_{4}$ and $v$ as in Definition 4.11. Let $\pi\in\left(\frac{1}{2},1\right)$ be a probability satisfying
@@
Suppose that Assumptions 2.1, 4.5 hold. Let $\bar{\mu}(\pi^{\frac{1}{4}})>0$ be defined as in Assumption 2.1(ii) and assume $\mu\leq\kappa\bar{\mu}(\pi^{\frac{1}{4}})$. Let $\psi_{k}$ be defined as in (4.13) and (4.14). Then, there exists $\sigma>0$ such that
@@
In order to obtain the expected value for the hitting time $\mathcal{K}_{\epsilon}$ given in Definition 4.1, we need the value $\delta^{\dagger}$, independent of $k$, such that if $\delta_{k}<\delta^{\dagger}$, iteration $k$ is true and $k<\mathcal{K}_{\epsilon}$, then the iteration is successful.
@@
(4.27) $\delta_{k}<\delta^{\dagger}:=\min\left\{\frac{\epsilon}{2\kappa},\frac{\epsilon}{2\eta_{2}},\frac{\underline{\theta}\epsilon(1-\eta_{1})}{2\theta_{0}(3\kappa+L)}\right\}.$
@@
Clearly, $W_{k}$ takes values ±$\pm 1$. Then, we can prove the following result.
@@
    there exists a nondecreasing function :[0,)($\ell:[0,\infty)\rightarrow(0,\infty)$ and a constant $\Theta>0$ such that, for all $k\geq 0$,
@@
and assume that $\delta_{\epsilon}=\gamma^{j_{\epsilon}}\delta_{0}$, for some integer $j_{\epsilon}\leq 0$; notice that we can always choose $\xi$ sufficiently large so that this is true. As a consequence, $\Delta_{k}=\gamma^{i_{k}}\delta_{\epsilon}$ for some integer $i_{k}$.
When $\mathbbm{1}_{\{\mathcal{K}_{\epsilon}>k\}}=0$, inequality (4.29) trivially holds. Otherwise, conditioning on $\mathbbm{1}_{\{\mathcal{K}_{\epsilon}>k\}}=1$, we can prove that
@@ 5. Numerical experiments
with $\xi_{i}$, $i=1,\ldots,m$, generated from a uniform random distribution on $[-\sigma,\sigma]$ for some $\sigma>0$. We consider 17 nonlinear least-squares problems taken from [20], which represent the overall set of problems extracted from References n. $8,21,26,36$ in [20, Section 2]. The selected problems are reported in Table 1. Note that the dimension of the unknown is set to $n=100$ for each problem.
@@
For each problem and algorithm, we performed $10$ runs using the level of noise $\sigma=10^{-1}$ as in [15, §6] and the initial guesses given in [20]. We adopt the number of samples employed in the average function and gradient approximations as the measure of the cost of the algorithms; more precisely, we use a counter Cost that is increased, at each function or gradient evaluation, by the number of samples employed at such evaluation. Furthermore, we impose a computational budget equal to $10^{5}(n+1)$ evaluations for irerm_v1 and storm_v1, and equal to $10^{4}(n+1)$ evaluations for irerm_v2 and storm_v2. We observe that irerm requires three function evaluations and one gradient evaluation, whereas storm asks for two function evaluations and one gradient evaluation. For all algorithms, we stop the iterations when either the maximum computational budget is exceeded or a maximum number of $k_{\max}=500$ iterations is met. Note that the upper bound $k_{\max}$ was never reached in our numerical tests.
@@
p1 Chained Rosenbrock function [20, Problem 2.1] 100 $2(n-1)$
p2 Chained Wood function [20, Problem 2.2] 100 $3(n-2)$
p3 Chained Powell singular function [20, Problem 2.3] 100 $2(n-2)$
p4 Chained Gragg and Levy function [20, Problem 2.4] 100 $25(n-2)/$ 2
@@
p7 Chained Freudstein and Roth function [20, Problem 2.7] 100 $2(n-1)$
p8 Toint quadratic merging problem [20, Problem 2.9] 100 $3(n-2)$
p9 Chained exponential problem [20, Problem 2.10] 100 $2(n-1)$
@@
p15 edensch [20, Problem 2.82] 100 $3(n-1)$
p16 genhumps [20, Problem 2.83] 100 $3(n-1)$
p17 errinros (modified) [20, Problem 2.84] 100 $2(n-1)$
@@ 6. Conclusions and future work
We proposed a trust-region algorithm with first-order random models suitable for the minimization of unconstrained noisy functions. The proposed algorithm, named irerm, is based on a constrained reformulation of the minimization problem typical of the Inexact Restoration approach [1, 12, 14, 17], where satisfying the constraint is equivalent to computing both the function and its gradient exactly. To our knowledge, this is the first attempt to adopt the Inexact Restoration approach for a general problem where the evaluation of both the objective function and its gradient is random and sufficiently accurate in probability, and to provide a theoretical analysis under conditions viable for the context. At each iteration, irerm first enforces the sufficient accuracy in probability of the function and gradient estimates, and then employs an acceptance test involving both the noisy function and the infeasibility measure. Based on the theoretical analysis in [9, 15], we derived the expected number of iterations needed to achieve an approximate first-order optimality point. A preliminary numerical illustration of irerm on a collection of nonlinear least-squares problems is also presented. As future work, we plan to deepen the numerical investigation of irerm as well as design a generalized version of irerm with second-order random models.
@@ 7. Appendix
This appendix is devoted to the proof of Theorem 4.12.
@@
The proof follows closely [15, Theorem 4.11]. Given $\psi_{k}$ in (4.13), the successive difference $\psi_{k+1}-\psi_{k}$ can be written as
@@
(7.1) $\quad+(1-v)(\delta_{k+1}^{2}-\delta_{k}^{2}),$
where the last inequality follows from (4.14) and the monotonicity of $\{\theta_{k}\}$ (see Lemma 4.2(i)). If $k$ is unsuccessful, then $x_{k+1}=x_{k}$, $y_{k+1}=y_{k}$, $\delta_{k+1}=\delta_{k}/\gamma$ and (7) reads
To proceed, we distinguish the case $\|\nabla f(x_{k})\|\geq\zeta\delta_{k}$ from the case $\|\nabla f(x_{k})\|<\zeta\delta_{k}$.
(i) Case $\|\nabla f(x_{k})\|\geq\zeta\delta_{k}$. We analyse separately the case of true and false iteration $k$. If the iteration $k$ is true, we can combine (4.5) with $\|\nabla f(x_{k})\|\geq\zeta\delta_{k}$ to obtain
where the last inequality is due to condition (4.21) and the upper bound $\mu\leq\kappa\bar{\mu}_{up}$, which follows from the hypothesis $\mu\leq\kappa\bar{\mu}(\pi^{\frac{1}{4}})$ and Assumption 2.1(ii). Since $k$ is true, $\kappa$ satisfies condition (4.20), and inequality (7.3) holds, Lemma 4.6 implies that $k$ is successful. Therefore, (3.3) and inequality (7) yield
@@
By summing and subtracting the quantities $\bar{f}_{k}^{*}$ and $\bar{f}_{k}^{p}$ in the above inequality, we get
@@
By the definition of $\operatorname{Ared}_{k}$ in (3.7), the fact that $y_{k+1}=y_{k+1}^{t}$ due to $k$ being successful, and (4.4)-(4.6) we get
@@
Since we know that (4.10) holds for successful iterations, from the previous inequality we easily get
We exploit once again the fact that $k$ is true and $\|\nabla f(x_{k})\|\geq\zeta\delta_{k}$, and observe that
@@
$-vC_{1}\delta_{k}\|\nabla f(x_{k})\|+(1-v)(\gamma^{2}-1)\delta_{k}^{2}\leq(-vC_{1}\zeta+(1-v)(\gamma^{2}-1))\delta_{k}^{2}<0.$
If the iteration $k$ is false, $k$ can be either a successful or an unsuccessful iteration. If $k$ is successful, then $x_{k+1}=x_{k}+p_{k}$, and applying the Descent Lemma for functions with Lipschitz continuous gradient to $f$ leads to the following chain of inequalities
$f(x_{k+1})-f(x_{k})$ $\leq\nabla f(x_{k})^{T}p_{k}+L\|p_{k}\|^{2}$
(7.7) $\leq\|\nabla f(x_{k})\|\delta_{k}+L\delta_{k}^{2}$
(7.8) $\leq\left(1+\frac{L}{\zeta}\right)\|\nabla f(x_{k})\|\delta_{k},$
where the last inequality follows from $\|\nabla f(x_{k})\|\geq\zeta\delta_{k}$. Furthermore, we have
@@
where the second inequality follows from $\|\nabla f(x_{k})\|\geq\zeta\delta_{k}$, the bounds $10<\theta_{k+1}<$ 1, the hypothesis $\mu\leq\kappa\bar{\mu}(\pi^{\frac{1}{4}})\leq\kappa\bar{\mu}_{up}$ and the definition of $C_{2}$ in (4.22).
@@
where the third inequality is again due to $\|\nabla f(X_{k})\|\geq\zeta\Delta_{k}$. Thus,
(7.13) $\mathbb{E}_{k-1}[\Psi_{k+1}-\Psi_{k}\,|\,\|\nabla f(X_{k})\|\geq\zeta\Delta_{k}]\leq-(1-v)(\gamma^{2}-1)\Delta_{k}^{2}.$
(ii) Case $\|\nabla f(x_{k})\|<\zeta\delta_{k}$. Again, we analyse true and false iterations separately. If the iteration $k$ is true and successful, then, reasoning as in Case (i), (7.4) still holds. Thus, (7.4) and (4.11) yield
@@
We observe that $02\kappa-\eta_{1}\eta_{2}<$ 0 by (4.20) and $\theta_{k+1}=\theta_{k+1}^{t}\geq\underline{\theta}$ due to the acceptance condition (3.10) holding for successful iterations. Then, by recalling the definition of $C_{3}$ in (4.22), we get
@@
If iteration $k$ is false and successful, then $x_{k+1}=x_{k}+p_{k}$, (7.7) gives
(7.15) $f(x_{k+1})-f(x_{k})$ $\leq\left(\zeta+L\right)\delta_{k}^{2},$
where the last inequality follows from $\|\nabla f(x_{k})\|<\zeta\delta_{k}$. Therefore, (7), (7.9) and (7.15) yield to
@@
Using (4.23), the choice of $\pi$ in (4.24) gives $(1-\pi)\left(vC_{4}+(1-v)\left(\frac{\gamma^{4}-1}{\gamma^{2}}\right)\right)\leq\frac{1}{2}(1-v)\left(\frac{\gamma^{2}-1}{\gamma^{2}}\right)$. Consequently,
(7.17) $\mathbb{E}_{k-1}[\Psi_{k+1}-\Psi_{k}\,|\,\|\nabla f(X_{k})\|<\zeta\Delta_{k}]\leq-\frac{1}{2}(1-v)\left(\frac{\gamma^{2}-1}{\gamma^{2}}\right)\Delta_{k}^{2}.$
Finally, by combining (7.13) and (7.17) and observing that $(\gamma^{2}-1)/(2\gamma^{2})<\gamma^{2}-1$, we obtain
(7.18) $\mathbb{E}_{k-1}[\Psi_{k+1}-\Psi_{k}]\leq-\frac{1}{2}(1-v)\left(\frac{\gamma^{2}-1}{\gamma^{2}}\right)\Delta_{k}^{2},$
and we conclude that (4.25) holds with $\sigma=\frac{1}{2}(1-v)(\gamma^{2}-1)/\gamma^{2}$.
In order to prove (4.26), summing (4.25) over $k=0,\ldots,K$, one gets
@@
By taking the total expected value and recalling the definition of conditional expected value, we obtain
$\mathbb{E}\left[\sum_{k=0}^{K}\sigma\Delta_{k}^{2}\right]$ $=\sum_{k=0}^{K}\mathbb{E}\left[\mathbb{E}_{k-1}[\Psi_{k}-\Psi_{k+1}]\right]$
@@
where the last inequality is due to the fact that $\{\Psi_{k}\}$ is bounded from below by a constant $\Psi_{low}$ for all $k\geq 0$, thanks to Assumption 2.1(i) and (iii). By taking the limit for $K\rightarrow\infty$ and applying the monotone convergence theorem on the left-hand side of the inequality, we get $\mathbb{E}\left[\sum_{k=0}^{\infty}\Delta_{k}^{2}\right]<\infty$ from which (4.26) immediately follows. ∎

### v3

@@ 4.3. Iteration complexity
We conclude our analysis by deriving the expected number $\mathbb{E}[{\mathcal{K}}_{\epsilon}]$ for the hitting time given in Definition 4.1. We rely on results for random processes introduced in [9, §2] and, to this end, we need some technical results that are proved in this section.
Let $v\in(0,1)$ be a prefixed constant to be later specified and $\Sigma$ a constant such that
(4.12) ${f(x_{k})}-h(y_{k})+\Sigma\geq 0,\quad\forall\ k\geq 0.$
We define the Lyapunov function
@@
(4.15) $\mathcal{G}_{k,1}$ $=\left\{|f(X_{k})-\bar{F}_{k}^{\dagger}|\leq\kappa\min\{\Delta_{k}^{2},h(Y_{k})\}\right\},$
(4.16) $\mathcal{G}_{k,2}$ $=\left\{|f(X_{k})-\bar{F}_{k}^{*}|\leq\kappa\min\{\Delta_{k}^{2},h(Y_{k})\}\right\},$
(4.17) $\mathcal{G}_{k,3}$ $=\left\{\|\nabla f(X_{k})-G_{k}\|\leq\kappa\Delta_{k}\right\},$
(4.18) $\mathcal{G}_{k,4}$ $=\left\{|f(X_{k}+P_{k})-\bar{F}_{k}^{p}|\leq\kappa\Delta_{k}^{2}\right\},$
@@
By Definition 4.3, we note that iteration $k$ is true when the realization of $\mathcal{I}_{k}\mathcal{J}_{k}$ equals 1.
@@
Note that Assumption 4.9 is satisfied whenever the random variables $\omega^{\dagger}_{y_{k+1}^{t}}$, $\omega^{*}_{y_{k+1}^{t}}$ and $\omega_{y_{k+1}^{g}}$ that model the noise are independent.
@@
Suppose that Assumptions 2.1 and 4.9 hold. Given a probability $\pi\in(0,1)$, let $\bar{\mu}(\pi^{\frac{1}{4}})>0$ be defined as in Assumption 2.1(ii) and assume that the parameter $\mu$ in Algorithm irerm is such that $\mu\leq\kappa\bar{\mu}(\pi^{\frac{1}{4}})$ with $\kappa$ as in Definition 4.3. Then we have
@@
Let $\kappa$, $\zeta$ be some scalars satisfying
@@
Letting $\Psi_{k}$ be the random function with realizations $\psi_{k}$, the following theorem shows a bound on the expected value $\mathbb{E}_{k-1}[\Psi_{k+1}-\Psi_{k}]$, from which one easily derives that $\Delta_{k}$ tends to zero almost surely.
@@
Let $\kappa,\,\zeta,\,C_{1},\,C_{2},\,C_{3},\,C_{4}$ and $v$ be as in Definition 4.11. Let $\pi\in\left(\frac{1}{2},1\right)$ be a probability satisfying
@@
Suppose that Assumptions 2.1, 4.5 and 4.9 hold. Let $\bar{\mu}(\pi^{\frac{1}{4}})>0$ be defined as in Assumption 2.1(ii) and assume $\mu\leq\kappa\bar{\mu}(\pi^{\frac{1}{4}})$ with $\kappa$ as in Definition 4.3. Let $\psi_{k}$ be defined as in (4.14) and (4.12). Then, there exists $\sigma>0$ such that
@@
In order to obtain the expected value for the hitting time $\mathcal{K}_{\epsilon}$ given in Definition 4.1, we need a value $\delta^{\dagger}$, independent of $k$, such that if $\delta_{k}<\delta^{\dagger}$, iteration $k$ is true and $k<\mathcal{K}_{\epsilon}$, then the iteration is successful. We specify the value of such $\delta^{\dagger}$ in the next Lemma.
@@
(4.27) $\delta_{k}<\delta^{\dagger}:=\min\left\{\frac{\epsilon}{2\kappa},\frac{\epsilon}{2\eta_{2}},\frac{\underline{\theta}\epsilon(1-\eta_{1})}{2(\theta_{0}(3\kappa+L)+\mu)}\right\}.$
@@
Clearly, $W_{k}$ takes values $\pm 1$. Then, we can prove the following result.
@@
    there exists a nondecreasing function $\ell:[0,\infty)\rightarrow(0,\infty)$ and a constant $\Theta>0$ such that, for all $k\geq 0$,
@@
and assume that $\delta_{\epsilon}=\gamma^{j_{\epsilon}}\delta_{0}$, for some integer $j_{\epsilon}\leq 0$; notice that we can always choose $\xi$ sufficiently large so that this is true. As a consequence, $\Delta_{k}=\gamma^{i_{k}}\delta_{\epsilon}$ for some random integer variable $i_{k}$.
When $\mathbbm{1}_{\{\mathcal{K}_{\epsilon}>k\}}=0$, inequality (4.29) trivially holds. Otherwise, conditioning on $\mathbbm{1}_{\{\mathcal{K}_{\epsilon}>k\}}=1$, let us show that
@@ 5. Numerical experiments
with $\xi_{i}$, $i=1,\ldots,m$, generated from a uniform random distribution on $\sigma,\sigma]$ for some $\sigma>0$. We consider 17 nonlinear least-squares problems taken from [20], which represent the overall set of problems extracted from References n. $8,21,26,36$ in [20, Section 2]. The selected problems are reported in Table 1. Note that the dimension of the unknown is set to $n=100$ for each problem.
@@
For each problem and algorithm, we performed $10$ runs using the level of noise $\sigma=10^{-1}$ as in [15, §6] and the initial guesses given in [20]. We adopt the number of samples employed in the average function and gradient approximations as the measure of the cost of the algorithms; more precisely, we use a counter Cost that is increased, at each function or gradient evaluation, by the number of samples employed at such evaluation. Furthermore, we impose a computational budget equal to $10^{5}(n+1)$ evaluations for irerm_v1 and storm_v1, and equal to $10^{4}(n+1)$ evaluations for irerm_v2 and storm_v2. We observe that irerm requires three function evaluations and one gradient evaluation, whereas storm asks for two function evaluations and one gradient evaluation per iteration. For all algorithms, we stop the iterations when either the maximum computational budget is exceeded or a maximum number of $k_{\max}=500$ iterations is met. Note that the upper bound $k_{\max}$ was never reached in our numerical tests.
@@
p1 Chained Rosenbrock function [20, Problem 2.1] 100 2(n1)2(n-1)
p2 Chained Wood function [20, Problem 2.2] 100 3(n2)3(n-2)
p3 Chained Powell singular function [20, Problem 2.3] 100 2(n2)2(n-2)
p4 Chained Gragg and Levy function [20, Problem 2.4] 100 5(n2)/25(n-2)/2
@@
p7 Chained Freudstein and Roth function [20, Problem 2.7] 100 2(n1)2(n-1)
p8 Toint quadratic merging problem [20, Problem 2.9] 100 3(n2)3(n-2)
p9 Chained exponential problem [20, Problem 2.10] 100 2(n1)2(n-1)
@@
p15 edensch [20, Problem 2.82] 100 3(n1)3(n-1)
p16 genhumps [20, Problem 2.83] 100 3(n1)3(n-1)
p17 errinros (modified) [20, Problem 2.84] 100 2(n1)2(n-1)
@@ 6. Conclusions and future work
We proposed a trust-region algorithm with first-order random models suitable for the minimization of unconstrained noisy functions. The proposed algorithm, named irerm, is based on a constrained reformulation of the minimization problem typical of the Inexact Restoration approach [1, 12, 14, 17], where satisfying the constraint is equivalent to computing both the function and its gradient exactly. To our knowledge, this is the first attempt to adopt the Inexact Restoration approach for a general problem where the evaluation of both the objective function and its gradient is random and sufficiently accurate in probability, and to provide a theoretical analysis under conditions viable for the context. At each iteration, irerm first enforces the sufficient accuracy in probability of the function and gradient estimates, and then employs an acceptance test involving both the noisy function and the infeasibility measure. Based on the theoretical analysis in [9, 15], we derived an upper bound on the expected number of iterations needed to achieve an approximate first-order optimality point. A preliminary numerical illustration of irerm on a collection of nonlinear least-squares problems is also presented. As future work, we plan to deepen the numerical investigation of irerm as well as design a generalized version of irerm with second-order random models.
@@ 7. Appendix
This appendix is devoted to the proof of Theorem 4.12, whose arguments follow closely those of [15, Theorem 4.11]. To this aim, we need the following four technical lemmas.
Lemma 7.1.
Suppose Assumption 4.5 holds and that iteration $k$ is successful. Then we have
(7.1) $f(x_{k+1})-f(x_{k})\leq\|\nabla f(x_{k})\|\delta_{k}+L\delta_{k}^{2}.$
@@
If the iteration is successful, then $x_{k+1}=x_{k}+p_{k}$. Applying the Descent Lemma for functions with Lipschitz continuous gradient to $f$ and using $p_{k}=-\delta_{k}g_{k}/\|g_{k}\|$ yields
$f(x_{k+1})-f(x_{k})\leq\nabla f(x_{k})^{T}p_{k}+L\|p_{k}\|^{2}\leq\|\nabla f(x_{k})\|\delta_{k}+L\delta_{k}^{2}.$
Lemma 7.2.
Suppose Assumption 2.1 holds and let $\psi_{k}$ be defined as in (4.13) and (4.14), with the constant $\Sigma$ chosen as in (4.12). The following facts hold true.
$(i)$ It holds
$\psi_{k+1}-\psi_{k}$ $\leq v\left(\theta_{k+1}({f(x_{k+1})}-{f(x_{k})})+(1-\theta_{k+1})(h(y_{k+1})-h(y_{k}))\right)$
(7.2) $\quad+(1-v)(\delta_{k+1}^{2}-\delta_{k}^{2}).$
$(ii)$ If iteration $k$ is true and successful, then
$(iii)$ If iteration $k$ is unsuccessful, then
Proof.
(i) The successive difference $\psi_{k+1}-\psi_{k}$ can be written as
@@
$\quad+(1-v)(\delta_{k+1}^{2}-\delta_{k}^{2}),$
where the last inequality follows from (4.12) and the monotonicity of $\{\theta_{k}\}$ (see Lemma 4.2(i)).
(ii) The occurrence of successful iteration (3.3) and inequality (7.2) yield
@@
By summing and subtracting the quantities $v\theta_{k+1}\bar{f}_{k}^{*}$ and $v\theta_{k+1}\bar{f}_{k}^{p}$ in the above inequality, we get
@@
By the definition of $\operatorname{Ared}_{k}$ in (3.7), the fact that $y_{k+1}=y_{k+1}^{t}$ due to $k$ being successful, and (4.5)-(4.7), we get
@@
Since we know that (4.11) holds for successful iterations, from the previous inequality we easily get (7.3).
(iii) If $k$ is unsuccessful, the claim follows from (7.2) and $x_{k+1}=x_{k}$, $y_{k+1}=y_{k}$, $\delta_{k+1}=\delta_{k}/\gamma$. ∎
To proceed, we distinguish the case $\|\nabla f(X_{k})\|\geq\zeta\Delta_{k}$ from the case $\|\nabla f(X_{k})\|<\zeta\Delta_{k}$, where $\zeta$ is defined in (4.21).
Lemma 7.3.
Under the Assumptions of Theorem 4.12, it holds
(7.6) $\mathbb{E}_{k-1}[\Psi_{k+1}-\Psi_{k}\,|\,\|\nabla f(X_{k})\|\geq\zeta\Delta_{k}]\leq-(1-v)(\gamma^{2}-1)\Delta_{k}^{2},$
with $\zeta$ as in (4.21).
Proof.
Let $x_{k}$, $\delta_{k}$ and $\psi_{k}$ denote the realizations of the random variables $X_{k},\Delta_{k}$ and $\Psi_{k}$, respectively. Consider an arbitrary realization of Algorithm irerm and the case $\|\nabla f(x_{k})\|\geq\zeta\delta_{k}$. We analyze separately true and false iterations. If the iteration $k$ is true, we can combine (4.6) with $\|\nabla f(x_{k})\|\geq\zeta\delta_{k}$ to obtain
where the last inequality is due to condition (4.21) and the upper bound $\mu\leq\kappa\bar{\mu}_{up}$, which follows from the hypothesis $\mu\leq\kappa\bar{\mu}(\pi^{\frac{1}{4}})$ and Assumption 2.1(ii). Since $k$ is true, $\kappa$ satisfies condition (4.20), and inequality (7.7) holds, Lemma 4.6 implies that $k$ is successful and consequently (7.3) holds.
By exploiting once again the fact that $k$ is true and $\|\nabla f(x_{k})\|\geq\zeta\delta_{k}$, we observe that
@@
$vC_{1}\delta_{k}\|\nabla f(x_{k})\|+(1-v)(\gamma^{2}-1)\delta_{k}^{2}\leq(-vC_{1}\zeta+(1-v)(\gamma^{2}-1))\delta_{k}^{2}<0.$
If the iteration $k$ is false, $k$ can be either a successful or an unsuccessful iteration. If $k$ is successful, using (7.1) and $\|\nabla f(x_{k})\|\geq\zeta\delta_{k}$ we obtain
(7.9) $f(x_{k+1})-f(x_{k})\leq\left(1+\frac{L}{\zeta}\right)\|\nabla f(x_{k})\|\delta_{k}.$
Furthermore, we have
@@
where the second inequality follows from $\|\nabla f(x_{k})\|\geq\zeta\delta_{k}$, the bounds $0<\theta_{k+1}<1$, the hypothesis $\mu\leq\kappa\bar{\mu}(\pi^{\frac{1}{4}})\leq\kappa\bar{\mu}_{up}$ and the definition of $C_{2}$ in (4.22).
@@
where the third inequality is again due to $\|\nabla f(X_{k})\|\geq\zeta\Delta_{k}$. ∎
Now we analyze the case $\|\nabla f(X_{k})\|<\zeta\Delta_{k}$, where $\zeta$ is again defined in (4.21).
Lemma 7.4.
Under the Assumptions of Theorem 4.12, it holds
(7.14) $\mathbb{E}_{k-1}[\Psi_{k+1}-\Psi_{k}\,|\,\|\nabla f(X_{k})\|<\zeta\Delta_{k}]\leq-\frac{1}{2}(1-v)\left(\frac{\gamma^{2}-1}{\gamma^{2}}\right)\Delta_{k}^{2},$
with $\zeta$ as in (4.21).
Proof.
Let $x_{k}$, $\delta_{k}$ and $\psi_{k}$ denote the realizations of the random variables $X_{k},\Delta_{k}$ and $\Psi_{k}$, respectively. Consider an arbitrary realization of Algorithm irerm and the case $\|\nabla f(x_{k})\|<\zeta\delta_{k}$. We analyze separately true and false iterations.
If the iteration $k$ is true and successful, then inequality (7.3) holds. Thus, (7.3) and condition (3.9) yield
@@
We observe that $2\kappa-\eta_{1}\eta_{2}<0$ by (4.20) and $\theta_{k+1}=\theta_{k+1}^{t}\geq\underline{\theta}$ due to the acceptance condition (3.10) holding for successful iterations. Then, by recalling the definition of $C_{3}$ in (4.22), we get
@@
If iteration $k$ is false and successful, (7.1) and $\|\nabla f(x_{k})\|<\zeta\delta_{k}$ give
(7.16) $f(x_{k+1})-f(x_{k})$ $\leq\left(\zeta+L\right)\delta_{k}^{2}.$
Therefore, (7.2), (7.10) and (7.16) yield to
@@
Using (4.23), the choice of $\pi$ in (4.24) gives $(1-\pi)\left(vC_{4}+(1-v)\left(\frac{\gamma^{4}-1}{\gamma^{2}}\right)\right)\leq\frac{1}{2}(1-v)\left(\frac{\gamma^{2}-1}{\gamma^{2}}\right)$ and the claim follows. ∎
We conclude by deriving the proof of Theorem 4.12.
Proof of Theorem 4.12.
By combining (7.6) and (7.14) and observing that $(\gamma^{2}-1)/(2\gamma^{2})<\gamma^{2}-1$, we obtain
$\mathbb{E}_{k-1}[\Psi_{k+1}-\Psi_{k}]\leq-\frac{1}{2}(1-v)\left(\frac{\gamma^{2}-1}{\gamma^{2}}\right)\Delta_{k}^{2},$
which means that inequality (4.25) holds with $\sigma=\frac{1}{2}(1-v)(\gamma^{2}-1)/\gamma^{2}$.
In order to prove (4.26), we sum (4.25) over $k=0,\ldots,K$ to obtain
@@
By taking the total expected value, using the linearity of expectation and the law of total expectation, we obtain
$\mathbb{E}\left[\sum_{k=0}^{K}\sigma\Delta_{k}^{2}\right]$ $\leq\sum_{k=0}^{K}\mathbb{E}\left[\mathbb{E}_{k-1}[\Psi_{k}-\Psi_{k+1}]\right]$
@@
where the last inequality is due to the fact that $\{\Psi_{k}\}$ is bounded from below by a constant $\Psi_{low}$ for all $k\geq 0$, thanks to Assumption 2.1(i) and (iii). By taking the limit for $K\rightarrow\infty$ and applying the monotone convergence theorem on the left-hand side of the inequality, we get $\mathbb{E}\left[\sum_{k=0}^{\infty}\Delta_{k}^{2}\right]<\infty$, from which (4.26) immediately follows. ∎